OpenStudy (anonymous):

On the graph of the equation x - 4y = -16, what is the value of the y-intercept?

OpenStudy (anonymous):

well, first change the equation to a y=mx+b. x-4y=-16 -4y=x-16 y=-1/4x+4 then, the y intercept would be 4 because b is the y intercept (y=mx+b)